# The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism

### v3

@@ Abstract
We investigate the baryonic Tully-Fisher relation in low surface brightness galaxies selected from the Arecibo Legacy Fast ALFA survey. We find that the $\rm HI$-bearing low surface brightness galaxies still follow the baryonic Tully-Fisher relation of typical late-type galaxies, with a slope of approximately 4 in the baryonic mass versus rotational velocity diagram on the logarithmic scale, i.e., $M_{\rm{b}}\propto v_{\rm{rot}}^{4}$. Our findings suggest that the matter distributions in low surface brightness galaxies may resemble that of general late-type galaxies, and hint that low surface brightness galaxies may not originate from dark matter halos of low densities or stronger/weaker feedback processes, but may emerge from dark matter halos with high spin values.
@@
pubyear: 2024pagerange: The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation MechanismThe Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism
@@ 1 Introduction
Low surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997). Many isolated LSBGs have been observed to exhibit higher HI fractions (e.g., Catinella et al., 2018; Schombert & McGaugh, 2021; Du et al., 2015; He et al., 2020). However, some studies (e.g., Martin et al., 2019; Galaz et al., 2002) alternatively found that some LSBGs exhibit the same or even smaller gas fractions compared to high surface brightness galaxies (HSBGs). LSBGs usually have low star formation rates (Wyder et al., 2009; Rong et al., 2020b), low metallicities (Kuzio de Naray et al., 2004; Schombert & McGaugh, 2021), and tend to appear in low-density environment (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
@@
The baryonic Tully-Fisher relation (BTFR) is a correlation between the rotational velocity, $v_{\rm rot}$, and the baryonic mass, $M_{\rm b}$, of galaxies. Over the past two decades, many studies have shown that late-type galaxies and dwarf galaxies consistently adhere to a power-law form of BTFR, despite their wide ranges of baryonic mass, luminosity, and size. BTFR could be written as, $M_{\rm b}\propto v_{\rm rot}^{\beta}$, where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g., McGaugh et al., 2000; Verheijen, 2001; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016a; Sales et al., 2017; Ponomareva et al., 2018; Goddy et al., 2023). Notably, some studies have also discovered that early-type galaxies may also follow the BTFR of late-type galaxies (den Heijer et al., 2015; Serra et al., 2012).
@@
As anticipated, recent studies examining the BTFR of LSBGs also show some potential signs of tension. Previous studies have found that the BTFR does not depend on galactic surface brightness or other galactic properties(e.g., Lelli et al., 2016a; Ponomareva et al., 2018; see also Zwaan et al., 1995). However, some recent investigations (e.g., Mancera Piña et al., 2019; Guo et al., 2020; Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that, the particular subsample of LSBGs with large disk sizes, i.e., UDGs, may significantly deviate from the BTFR of typical late-type galaxies, indicative of the possible dependence on size (Mancera Piña et al., 2020). Therefore, a thorough and meticulous examination of the BTFR for LSBGs becomes essential. Such an analysis may not only indicate whether the matter distribution within LSBGs aligns with that of typical late-type and dwarf galaxies, but also provide valuable constraints on the formation mechanisms of LSBGs.
@@ 4 BTFR
Our results are consistent with many previous observational (e.g., McGaugh, 2012; Lelli et al., 2016a; Papastergis et al., 2016) and simulation (e.g., Dutton et al., 2017) results. However, it is also worth noting that some other studies alternatively report smaller slopes ($\beta\sim 3$, e.g., Goddy et al., 2023; Ponomareva et al., 2018).
Figure 3: The $v_{\rm rot}-M_{\rm b}$ diagram for our samples. The green- and the grey-filled circles are sthe same as those in Figure 2. The orange-filled circles are SPARC galaxies from Lelli et al. (2019). All the error bars denote $\rm 1\sigma$ uncertainties. The blue squares represent the median values of LSBGs while the black squares represent those of HSBGs. The red-solid line is the tightest BTFR reported by Lelli et al. (2019). The red-shaded area is the $\rm 1\sigma$ intrinsic scatter of the red-solid line, which is considered as the $\rm 1\sigma$ uncertainty of the tightest BTFR.
@@ 5.1 Potential biases
Additionally, possible bias emerges from using $W_{20}$ or $W_{50}$ (the $50\%$ peak width of an HI spectrum) to estimate the rotation velocities. However, we find that the average difference between the $\log v_{\rm rot}$ estimated from $W_{20}$ and $W_{50}$ is quite small, even smaller than the uncertainties of $\log v_{\rm rot}$. Therefore, using $W_{20}$ or $W_{50}$ would not significantly affect our results. Besides, we also test whether using the different methods of estimating $M_{\rm*}$ (Taylor et al., 2011) or considering the intrinsic thickness $q_{0}$ dependence on $M_{\rm*}$ (Lelli et al., 2016b) would lead to the changes of BTFR slopes of LSBGs. We observe no strong changes in BTFR slopes in these tests. Furthermore, we notice that the surface brightness limit determined by the optical survey may lead to large uncertainties in $b/a$ estimation for LSBGs (Mancera Piña et al., 2024). However, the large uncertainties should result in a larger scatter, rather than a systematic bias or a change of BTFR slope.

### v4

# The baryonic Tully-Fisher relation of HI-bearing low-surface brightness galaxies implies their formation mechanism

@@ Abstract
We investigate the baryonic Tully-Fisher relation in low-surface brightness galaxies selected from the Arecibo Legacy Fast ALFA survey. We find that the $\rm HI$-bearing low-surface brightness galaxies still follow the baryonic Tully-Fisher relation of typical late-type galaxies, with a slope of approximately 4 in the baryonic mass versus rotational velocity diagram on the logarithmic scale, i.e., $M_{\rm{b}}\propto v_{\rm{rot}}^{4}$. Our findings suggest that the matter distributions in low-surface brightness galaxies may resemble that of general late-type galaxies, and hint that low-surface brightness galaxies may not originate from dark matter halos of low densities or stronger/weaker feedback processes, but may emerge from dark matter halos with high spin values.
@@
pubyear: 2024pagerange: The baryonic Tully-Fisher relation of HI-bearing low-surface brightness galaxies implies their formation mechanismThe baryonic Tully-Fisher relation of HI-bearing low-surface brightness galaxies implies their formation mechanism
@@ 1 Introduction
low-surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997). Many isolated LSBGs have been observed to exhibit higher HI fractions (e.g., Catinella et al., 2018; Schombert & McGaugh, 2021; Du et al., 2015; He et al., 2020). However, some studies (e.g., Martin et al., 2019; Galaz et al., 2002) alternatively found that some LSBGs exhibit the same or even smaller gas fractions compared to high-surface brightness galaxies (HSBGs). LSBGs usually have low star formation rates (Wyder et al., 2009; Rong et al., 2020b), low metallicities (Kuzio de Naray et al., 2004; Schombert & McGaugh, 2021), and tend to appear in low-density environment (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
@@
The baryonic Tully-Fisher relation (BTFR) is a correlation between the rotational velocity, $v_{\rm rot}$, and the baryonic mass, $M_{\rm b}$, of galaxies. Over the past two decades, many studies have shown that late-type galaxies and dwarf galaxies consistently adhere to a power-law form of BTFR, despite their wide ranges of baryonic mass, luminosity, and size. BTFR could be written as, $M_{\rm b}\propto v_{\rm rot}^{\beta}$, where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g., McGaugh et al., 2000; Verheijen, 2001; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016b; Sales et al., 2017; Ponomareva et al., 2018; Goddy et al., 2023). Notably, some studies have also discovered that early-type galaxies may also follow the BTFR of late-type galaxies (den Heijer et al., 2015; Serra et al., 2012).
@@
As anticipated, recent studies examining the BTFR of LSBGs also show some potential signs of tension. Previous studies have found that the BTFR does not depend on galactic surface brightness or other galactic properties(e.g., Lelli et al., 2016b; Ponomareva et al., 2018; see also Zwaan et al., 1995). However, some recent investigations (e.g., Mancera Piña et al., 2019; Guo et al., 2020; Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that, the particular subsample of LSBGs with large disk sizes, i.e., UDGs, may significantly deviate from the BTFR of typical late-type galaxies, indicative of the possible dependence on size (Mancera Piña et al., 2020). Therefore, a thorough and meticulous examination of the BTFR for LSBGs becomes essential. Such an analysis may not only indicate whether the matter distribution within LSBGs aligns with that of typical late-type and dwarf galaxies, but also provide valuable constraints on the formation mechanisms of LSBGs.
@@ 4 BTFR
Our results are consistent with many previous observational (e.g., McGaugh, 2012; Lelli et al., 2016b; Papastergis et al., 2016) and simulation (e.g., Dutton et al., 2017) results. However, it is also worth noting that some other studies alternatively report smaller slopes ($\beta\sim 3$, e.g., Goddy et al., 2023; Ponomareva et al., 2018).
Figure 3: The $v_{\rm rot}-M_{\rm b}$ diagram for our samples. The green- and the grey-filled circles are sthe same as those in Figure 2. The orange-filled circles are SPARC galaxies (Lelli et al., 2016a) from Lelli et al. (2019). All the error bars denote $\rm 1\sigma$ uncertainties. The blue squares represent the median values of LSBGs while the black squares represent those of HSBGs. The red-solid line is the tightest BTFR reported by Lelli et al. (2019). The red-shaded area is the $\rm 1\sigma$ intrinsic scatter of the red-solid line, which is considered as the $\rm 1\sigma$ uncertainty of the tightest BTFR.
@@ 5.1 Potential biases
Additionally, possible bias emerges from using $W_{20}$ or $W_{50}$ (the $50\%$ peak width of an HI spectrum) to estimate the rotation velocities. However, we find that the average difference between the $\log v_{\rm rot}$ estimated from $W_{20}$ and $W_{50}$ is quite small, even smaller than the uncertainties of $\log v_{\rm rot}$. Therefore, using $W_{20}$ or $W_{50}$ would not significantly affect our results. Besides, we also test whether using the different methods of estimating $M_{\rm*}$ (Taylor et al., 2011) or considering the intrinsic thickness $q_{0}$ dependence on $M_{\rm*}$ (Lelli et al., 2016c) would lead to the changes of BTFR slopes of LSBGs. We observe no strong changes in BTFR slopes in these tests. Furthermore, we notice that the surface brightness limit determined by the optical survey may lead to large uncertainties in $b/a$ estimation for LSBGs (Mancera Piña et al., 2024). However, the large uncertainties should result in a larger scatter, rather than a systematic bias or a change of BTFR slope.